OpenStudy (anonymous):

I have this exercise and i don't now how to do it... a= -2/3 b= +7/3 I have do do this : -a-b /-a+b Please expalin me how to do it !!! :)

OpenStudy (anonymous):

(2/3 -7/3) / (2/3 + 7/3) = (-5/3) / 7/3 (-5/3) / (7/3) = (-5/3)(3/9) = -15/27

OpenStudy (anonymous):

i'm not sure but i think that best way to solve